# Deep Unlearn: Benchmarking Machine Unlearning

### v1

@@ 12 FF 0 0 0 9 12 FF 0 0 0 9
Figure 1: U-LiRA on CIFAR-10 on ResNet models. Both CT and MSG, which ranked first against U-MIA, showed great resilience against the U-LiRA attack.
@@ 6 Discussion and Conclusion
The increasing focus on data privacy and trustworthiness of machine learning models underscores the need for robust and practical methods to unlearn and remove the influence of specific data from trained models. Due to the growing size of models, we require methods that avoid the computationally costly retraining from scratch. In this work we performed a comprehensive comparison of approximate unlearning methods across various models and datasets aimed to address this critical issue.
@@
We experimentally compared 18 methods across different datasets and architectures, focusing on assessing the method’s ability to maintain privacy and accuracy while being computationally efficient and reliable across datasets, architectures and random seeds. Our findings indicate that Masked-Small-Gradients, which accumulates gradients via gradient descent on the data to remember and gradient ascent on the data to forget to determine which weights to update, consistently outperforms for all metrics across the studied datasets, architectures, and initialization seeds. Similarly, Convolution Transpose, which leverages the simple transposition in convolutional layers, performed strongly.
@@
Both CT and MSG were resistant against both a population-based Membership Inference Attack (MIA) and a stronger, per-sample attack (U-LiRA). However, a core challenge of approximate unlearning is that these methods will only be as strong as the attacks against which they are tested. As stronger and more complex attacks emerge, some approximate unlearning methods might no longer be as efficient as initially expected. This highlights the need for continuous evaluation and adaptation of unlearning methods to maintain their effectiveness. We also conducted experiments based on L2 distances, but found that no method consistently got close to the reference models’ weights, we provide further information in Appendix G.
@@
Limitations. Due to computational costs, we limited our analysis to Tiny Vision Transformers and ResNet; a further investigation of other architectures could provide useful insights. We did not investigate different amounts of unlearning samples, which some methods are known to be sensitive to [37]. We did not consider repeated deletion, instead we assume that there is a single forget set and that the unlearning process happens once, as is common in the literature, nonetheless, in practical applications one might need to unlearn different smaller forget sets over time and some unlearning methods might not work as well under such scenario. We finally remark once again on the difficulty of evaluation for approximate unlearning [31]: while these methods provide significant gains in efficiency, novel attacks might highlight yet unknown weaknesses of the unlearning processes.
@@
Future work. First, we put our focus on natural image data, however, machine unlearning is relevant to other data types such as medical images or other modalities such as time series, audio and speech, or language data. Second, we focus on the classification task, however, other learning tasks would greatly benefit from machine unlearning too. For instance removing concepts from generative models for images [20] or poisoned data in language models [31]. Third, this work focuses on empirically benchmarking approximate machine unlearning methods. We do not provide a theoretical analysis of these methods or a rigorous comparison with exact unlearning algorithms.
@@
Impact statement. This paper aims to highlight the importance of effectively assessing approximate machine unlearning methods. Our goal is to stress the need for evaluating new unlearning methods against more reliable baselines and experimental setups. Additionally, it is crucial to assess the consistency of a new unlearning method across various datasets and model architectures. Without such a thorough evaluations, proposed unlearning methods may provide a false sense of privacy and safety, ultimately limiting their effectiveness for data regulation.
7 Acknowledgements
@@
Xavier F. Cadet is supported by UK Research and Innovation (UKRI Centre for Doctoral Training in AI for Healthcare grant number EP/S023283/1). Hamed Haddadi is supported by the EPSRC Open Plus Fellowship (EP/W005271/1: Securing the Next Billion Consumer Devices on the Edge).
8 Impact Statement
This paper presents work whose goal is to advance the fields of Machine Learning and Machine Unlearning. There are many potential societal consequences of our work. We aim to raise awareness on the necessity of developing robust methods and evaluating the consistency of Machine Unlearning methods across datasets, as relying on inconsistent Machine Unlearning mechanism could go against Data Regulations.
@@ 8 Impact Statement
  • Related Works (Section A) provides an overview of the previous works related to benchmarking MU.
  • Datasets (Section B) contains information about the datasets, data augmentations, and hyper-parameters used to train the Original and Retrained models.
  • Neural Network Architectures (Section C) provides information about ResNet18 and TinyViT used throughout our benchmark.
  • Privacy Evaluation (Section D) gives the descriptions of the U-MIA and U-LiRA evaluation metrics.
@@
  • Per Dataset Results (Section E) shows some experimental results in terms of accuracy, retention, privacy metrics, and runtime efficiency per dataset for the 9 combinations of datasets and DNN architectures considered in our work.
@@
  • Per Architectures Ranking (Section F) provides Performance Retention Deviation and Indiscernibility Rankings separated for ResNet18 and TinyViT.
@@
  • L2 Distances between Model Weights (Section G) shows L2 distances computed between the Unlearned, Original, and Retrained models.
  • Requirements (Section H) which describes the compute resources.
Appendix A Related Works
A.1 Machine Unlearning
Machine Unlearning is often first associated with the work from Cao et al. [11], followed by Bourtoule et al. [9], which proposes SISA (Sharded, Isolated, Sliced, Aggregated) as an exact unlearning method. For a recent overview of MU, we refer to the survey from Xu et al. [50], which provides a taxonomy of common unlearning methods. Furthermore, Zhang et al. [52] review MU through privacy-preserving and security lenses. The authors cover the Confidentiality, Integrity, and Availability security triad and the need for Data Lineage, which relates to following the movement of data in a machine learning pipeline and understand from where it originates, where it is stored, and how it percolates in the system through transformation. Some might have information on common MU verification methods, privacy evaluation metrics, and datasets, we defer to the work of Nguyen et al. [39], and Shaik et al. [43].
(1) Data Reorganization methods focus on directly modifying the data to perform unlearning. It is divided into Data Obfuscation, Data Pruning, and Data Replacement.
Data Pruning usually relies on dividing the dataset into multiple sub-datasets and training sub-models on these subsets. This is the category to which SISA [9] relates. Our work does not consider methods associated with this setting as they assume the training process.
Data Replacement attempts to unlearn by replacing the original dataset with transformed data that simplifies unlearning specific samples. For instance, Cao et al. [11], replace the training data with summations of efficiently computable transformations. Like data pruning, these methods tend to make strong assumptions about the training process.
(2) Model Manipulation methods directly adjust the model parameters to remove the influence of specific data points. Model manipulation is divided into Model Shifting, Model Replacement, and Model Pruning.
Model Replacement uses pre-calculated parameters that do not reflect the data to forget to replace parts of the trained model. For instance, when using decision trees, one can replace nodes affected by the forget set by pre-calculated node [41].
A.2 Machine Unlearning for Deep Neural Networks
Initially, MU research primarily focused on linear models such as linear regression and logistic models. Such models allow for the design of methods that assume the convexity of the loss function, rendering them less practical for DNN-based approaches. Since DNNs can memorize parts of their training data, they are particularly relevant targets for MU, even more so when they have been trained on large amounts of potentially personal data. For Deep Learning models, unlearning raises additional challenges: 1) the non-convexity of the loss function of Deep Neural Networks [15], 2) the size of the models inducing large computational costs, 3) the randomness coming from the model’s training process, such as the initialization seed, randomness in the mini-batch generation process, and 4) the fact that any model update impacts subsequent versions of the models, namely the weights at epoch $n+1$ directly depend on the weights at update $n$.
A.3 Post-Hoc Machine Unlearning
Thus, we focus on post-hoc MU, a scenario where we assume that the unlearning method is agnostic to the original training process of the model. Under such a scenario, differences exist in terms of data availability at unlearning time. For instance, whether one has access to the original training data $\mathcal{D}$, the retain set $\mathcal{D}_{R}$, the forget set $\mathcal{D}_{F}$, or even some external set such as the validation set $\mathcal{D}_{V}$. Therefore, careful consideration should be given to the data requirement associated with an unlearning method. Indeed, some might require having access to both $\mathcal{D}_{R}$ and $\mathcal{D}_{F}$ at the unlearning time, while others assume that $\mathcal{D}_{R}$ is no longer available [17] making them more practical in real-world scenarios. Throughout our benchmark, we make the same assumption as the NeurIPS2023 Unlearning Challenge [45], where the unlearning methods had access to $f_{O},\mathcal{D}_{R},\mathcal{D}_{F},\mathcal{D}_{V}$
A.4 Machine Unlearning and Differential Privacy
We based our Unlearning definition on Sekhari et al. [42] and refer to their work on the distinction between Differential Privacy and the objective of Machine Unlearning. Differential privacy, in a high-level picture, is a method for publicly sharing aggregated information about a population by describing the patterns discovered among the groups within the dataset while withholding specific information about individual data points. A randomized algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-differentially private if for all datasets $D_{1}$ and $D_{2}$ that differ on a single data point, and all $S\subseteq\text{Range}(\mathcal{A})$,
Appendix B Datasets
CIFAR-10
CIFAR-10 is a widely used dataset in computer vision and machine learning. It comprises 60,000 32x32 color images in 10 different classes, with 6,000 images per class. The dataset is divided into 50,000 training images and 10,000 testing images. CIFAR-10 represents a diverse range of everyday objects, such as airplanes, automobiles, birds, and cats, making it a challenging task for image classification. The simplicity of the images combined with the variety of categories makes CIFAR-10 a suitable dataset to test the efficacy of machine unlearning algorithms in effectively unlearning information without compromising the model’s performance on the remaining data.
CIFAR-100
CIFAR-100 is a more complex extension of CIFAR-10, containing 100 classes with 600 images per class, split into 500 training images and 100 testing images per class. Each class is labeled with a "fine" label and grouped into 20 "coarse" labels, adding another layer of classification difficulty. The increased number of classes and finer granularity make CIFAR-100 an intriguing dataset for machine unlearning benchmarks. It poses a more significant challenge for models to forget specific classes or groups while retaining knowledge of others, thus testing the unlearning algorithms’ precision and effectiveness in handling more granular and complex datasets.
MNIST
The MNIST dataset is a well-known benchmark in handwritten digit recognition. It comprises 70,000 grayscale images of handwritten digits (0-9), 60,000 used for training, and 10,000 for testing. Each image is 28x28 pixels in size. We consider MNIST due to its simplicity and extensive research and development history. The simplicity of MNIST allows researchers to focus on the fundamental aspects of unlearning techniques without the additional complexity introduced by color or high resolution, providing a clear assessment of the effectiveness of unlearning algorithms in a controlled setting.
Fashion MNIST
Fashion MNIST is a more challenging replacement for MNIST. It contains 70,000 grayscale images of fashion items in 10 categories: shirts, trousers, and sneakers. Like MNIST, each image is 28x28 pixels, but the increased complexity and variability of clothing items make it a more challenging classification task. Fashion MNIST provides a more realistic and intricate dataset than MNIST, testing the unlearning algorithms’ ability to handle real-world-like variability and ensuring that they can effectively remove learned information while maintaining performance on a moderately complex dataset.
UTKFace
For each dataset, the Original and Retrained models are trained using the same hyper-parameters (provided in Table 3)
@@ Appendix B Datasets
Table 3: Summary of the number epochs, learning rate, and batch size for each dataset and model used to train the Original and Retrained models.
@@
Appendix C Neural Network Architectures
ResNet: ResNet18
ViT: TinyViT
Vision Transformer (ViT), introduced by Dosovitskiy et al. [19], adapts the transformer architecture to image classification by treating images as sequences of patches. We consider TinyViT from Wu et al. [48], as it is a compact version of ViT designed to be parameter-efficient while maintaining high performance.
Appendix D Privacy Evaluation
D.1 Unlearning-Membership Inference Attack (U-MIA)
D.2 Unlearning -Likelihood Ratio Attack (U-LiRA)
The performance of a general MIA can be improved by considering, e.g., a per-sample attack such as LiRA [13, 31]. For any given point, we wish to determine whether the outputs from the unlearned models differ from those of models that have never seen the data point. To assess the attack robustly, we evaluate it across multiple models, using shadow models trained on various retain/forget sets. Specifically, we first train $n$ models based on $n$ splits of the training data. This train data is then split into 10 random retain and forget splits, and hence, we unlearn a total of $10n$ models. We then perform hyper-parameter sweeps, similar to what we do in the original results and unlearn using the optimal hyper-parameters, except that we consider $\frac{n}{2}$ sweeps and conduct $200$ trials per sweep to determine the best hyper-parameters. In our setting, we set $n=64$.
Appendix E Per dataset results
Here, we present the results for both ResNet18 and the TinyViT across datasets.
ResNet18 We provide the tables with Retain Accuracy (RA), Forget Accuracy (FA), Test Accuracy (TA), Retain Retention (RR), Forget Retention (FR), Test Retention (TR), Performance Retention Deviation (RetDev), Indiscinerbility concerning the Test Set (Indisc), U-MIA on the Test set (T-MIA) and RunTime Efficiency (RTE) for every dataset using the ResNet18 model on MNIST (Table 4), FashionMNIST (Table 5), CIFAR-10 (Table 6), CIFAR-100 (Table 7) and UTKFace (Table 8). In general, CIFAR-100 provides the most visible differences, as the performance on the retain set is much higher than on the test. Datasets such as MNIST and FashionMNIST tend to show smaller differences between the methods as the performance on both the Retain and Test sets are similar, to begin with.
Table 4: MNIST - ResNet18
@@ Appendix E Per dataset results
Table 5: FashionMNIST - ResNet18
@@
Table 6: CIFAR-10 - ResNet18
@@
Table 7: CIFAR-100 - ResNet18. CIFAR-100 provides the most visible comparison as there is a large gap in performance between the Retain Set and Test set, this leads to much larger RetDev scores.
RA FA TA RR FR TR RetDev Indisc T-MIA RTE
Unlearner
BT 0.98 0.68 0.54 0.98 1.25 0.99 0.27 0.95 0.48 9.39
CF-k 1.00 0.83 0.56 1.00 1.53 1.02 0.55 0.73 0.63 5.91
CFW 0.98 0.43 0.43 0.98 0.79 0.78 0.44 1.00 0.50 6.17
CT 0.99 0.53 0.53 0.99 0.97 0.97 0.07 0.99 0.49 11.82
EU-k - - - - - - - - - -
FCS 0.98 0.54 0.55 0.98 0.99 1.01 0.04 0.92 0.54 3.02
FF - - - - - - - - - -
FT 0.98 0.55 0.54 0.98 1.02 0.98 0.05 0.99 0.50 5.16
GA 0.34 0.33 0.24 0.34 0.60 0.44 1.61 0.90 0.55 39.97
IU - - - - - - - - - -
KDE 0.99 0.52 0.51 0.99 0.95 0.94 0.11 0.99 0.50 3.98
MSG 0.91 0.38 0.38 0.91 0.69 0.69 0.71 1.00 0.50 4.49
NG+ 0.89 0.59 0.49 0.89 1.08 0.89 0.29 0.98 0.49 12.14
O 0.98 0.98 0.56 0.98 1.81 1.02 0.85 0.53 0.73 1.10
PRMQ 0.97 0.47 0.46 0.97 0.86 0.85 0.32 1.00 0.50 4.34
R 1.00 0.55 0.55 1.00 1.00 1.00 0.00 0.99 0.49 1.00
RNI 0.99 0.45 0.45 0.99 0.83 0.82 0.36 0.98 0.49 3.65
SCRUB 0.97 0.50 0.53 0.97 0.91 0.96 0.15 0.98 0.51 3.81
SRL 1.00 0.55 0.52 1.00 1.00 0.95 0.06 0.98 0.49 3.67
SalUN 0.98 0.49 0.51 0.98 0.91 0.93 0.18 0.99 0.49 10.66
Table 8: UTKFace - ResNet18
@@
TinyViT We provide the tables with RA, FA, TA, RR, FR, TR, RetDev, Indisc, T-MIA and RTE for MNIST (Table 9), FashionMNIST (Table 10), CIFAR-10 (Table 11) and CIFAR-100 (Table 12) using the TinyViT model.
Table 9: MNIST - TinyViT
@@
Table 10: FashionMNIST - TinyViT
@@
Table 11: CIFAR-10 - TinyViT
@@
Table 12: CIFAR-100 - TinyViT
@@
Appendix F Per architectures rankings
Here, we present the rankings across datasets for ResNet18 (Table 13) and TinyVit (Table 14). We note that some methods, such as RNI or NG+, are less efficient on the ViT architectures regarding Indiscernibility. However, methods such as SCRUB are less efficient regarding Retention Deviation on the ViT architecture.
Table 13: Ranking on ResNet

### v2

@@ 6 Main Results
Figure 3: U-LiRA on CIFAR-10 on ResNet models. CT and MSG, which ranked first against U-MIA, showed great resilience against the U-LiRA attack.
6.6 On the performance when evaluated against a stronger MIA.
From the results above, we note that MU methods exist that perform fast and reliably across datasets, architectures, and initializations. Specifically, CT, FT, and MSG seem strong candidates for reliable MU methods. However, as has also been highlighted in prior work [62], MIA has been debated as a strong metric, as its ability to assess MU is hampered by its ability to infer data membership. In line with this, recent works have introduced more powerful variations on MIA [63] with [14] proposing the stronger U-LiRA attack for MU. For the best performers in Table III, we apply the attack setup from [45] where we generate a total of 640 models (with varying train, retain, and forget sets) and for each unlearn method, perform a hyperparameter sweep to find the best configuration (for details, see Appendix 5.5). We determine for each data point its U-LiRA Inference Accuracy and report for each method its average and standard deviation (see Figure 3). From this, we conclude that MSG, as well as CT, resist U-LiRA attacks. Both MSG and CT thus not only rank first in terms of Performance Retention and Indiscernibility based on U-MIA, but also are robust against a stronger variation of MIA.
@@ 7.1 Conclusion
The increasing focus on data privacy and the trustworthiness of machine learning models underscores the need for robust and practical methods to unlearn and remove the influence of specific data from trained models. Due to the growing size of models, we require methods that avoid computationally costly retraining from scratch. In this work, we comprehensively compared approximate unlearning methods across various models and datasets to address this critical issue.
@@
We compared 18 methods across different datasets and architectures, focusing on assessing the method’s ability to maintain privacy and accuracy while being computationally efficient and reliable across datasets, architectures, and random seeds. Our findings indicate that Masked-Small-Gradients, which accumulates gradients via gradient descent on the data to remember and gradient ascent on the data to forget to determine which weights to update, consistently outperforms all metrics across the datasets, architectures, and initialization seeds. Similarly, Convolution Transpose, which leverages the simple transposition in convolutional layers, performs well.
@@
CT and MSG were resistant against a population-based Membership Inference Attack (MIA) and a stronger, per-sample attack (U-LiRA). However, a core challenge of approximate unlearning is that these methods will only be as strong as the attacks against which they are tested. As stronger and more complex attacks emerge, some approximate unlearning methods might no longer be as efficient as initially expected. This highlights the need for continuous evaluation and adaptation of unlearning methods to maintain effectiveness. We also conducted experiments based on L2 distances but found that no method consistently got close to the reference models’ weights (see Appendix C).
7.2 Limitations
@@
Due to computational costs, we limited our analysis to Tiny Vision Transformers and ResNet; further investigating other architectures could provide useful insights. We did not investigate different amounts of unlearning samples, which some methods are known to be sensitive to [14]. We did not consider repeated deletion. Instead, we assume that there is a single forget set and that the unlearning process happens once, as is common in the literature. Nonetheless, in practical applications, one might need to unlearn different smaller forget sets over time, and some unlearning methods might not work as well under such a scenario. We finally remark once again on the difficulty of evaluation for approximate unlearning [45]: while these methods provide significant gains in efficiency, novel attacks might highlight yet unknown weaknesses of the unlearning processes. This paper focuses exclusively on vision data modalities using vision architectures for image classification. While computer vision is a significant application area, exploring other modalities and architectures to generalize our findings is important. Different data types and models may rely on distinct inductive biases, and extrapolating our results to these other domains presents challenges. For example, in audio and sequential data, the interdependence and contextual relationships between data points introduce additional complexities, making unlearning more difficult. Additionally, this study was conducted with a fixed unlearning budget of 10% of the data. While it would be valuable to analyze other unlearning budgets, we prioritized thoroughly investigating this particular budget, which is already commonly used in the community. Our analysis also focused on unlearning randomly selected data points rather than specific classes, features, or particularly challenging data. As a result, our findings may not fully generalize to these other scenarios. Nevertheless, we believe this setting can effectively represent actual deletion requests from users of an image classification service provider.
7.3 Future work
@@
First, we focus on natural image data; however, machine unlearning is relevant to other data types, such as medical images or other modalities, such as time series, audio, and speech, or language data. Second, we focus on the classification task; however, machine unlearning would greatly benefit other learning tasks as well. For instance removing concepts from generative models for images [12] or poisoned data in language models [45]. Third, this work focuses on empirically benchmarking approximate machine unlearning methods. We do not provide a theoretical analysis of these methods or a rigorous comparison with exact unlearning algorithms.
7.4 Impact statement
@@
This paper aims to highlight the importance of effectively assessing approximate machine unlearning methods. Our goal is to stress the need to evaluate new unlearning methods against more reliable baselines and experimental setups. Additionally, it is crucial to assess the consistency of a new unlearning method across various datasets, model architectures, and model initializations. Without such a thorough evaluation, proposed unlearning methods may provide a false sense of privacy and safety, ultimately limiting their effectiveness for data regulation.
Acknowledgment
@@
Xavier F. Cadet is supported by UK Research and Innovation (UKRI Centre for Doctoral Training in AI for Healthcare grant number EP/S023283/1). Hamed Haddadi is supported by the EPSRC Open Plus Fellowship (EP/W005271/1: Securing the Next Billion Consumer Devices on the Edge). For open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.
@@
  • Per Dataset Results (Section A) shows some experimental results in terms of accuracy, retention, privacy metrics, and runtime efficiency per dataset for the 9 combinations of datasets and DNN architectures considered in our work.
@@
  • Per Architectures Ranking (Section B) provides Performance Retention Deviation and Indiscernibility Rankings separated for ResNet18 and TinyViT.
@@
  • L2 Distances between Model Weights (Section C) shows L2 distances computed between the Unlearned, Original, and Retrained models.
  • Data Availability (Section D) describes the data availability.
  • Requirements (Section E) which describes the compute resources.
@@
TABLE VII: Summary of the number epochs, learning rate, and batch size for each dataset and model used to train the Original and Retrained models.
@@
Appendix A Per dataset results
Here, we present the ResNet18 and the TinyViT results across datasets.
ResNet18 We provide the tables with Retain Accuracy (RA), Forget Accuracy (FA), Test Accuracy (TA), Retain Retention (RR), Forget Retention (FR), Test Retention (TR), Performance Retention Deviation (RetDev), Indiscinerbility concerning the Test Set (Indisc), U-MIA on the Test set (T-MIA) and RunTime Efficiency (RTE) for every dataset using the ResNet18 model on MNIST (Table VIII), FashionMNIST (Table IX), CIFAR-10 (Table X), CIFAR-100 (Table IV) and UTKFace (Table XI). In general, CIFAR-100 provides the most visible differences, as the performance on the retain set is much higher than on the test. Datasets such as MNIST and FashionMNIST tend to show smaller differences between the methods as the performance on both the Retain and Test sets are similar, to begin with.
TABLE VIII: MNIST - ResNet18
@@ Appendix A Per dataset results
TABLE IX: FashionMNIST - ResNet18
@@
TABLE X: CIFAR-10 - ResNet18
@@
TABLE XI: UTKFace - ResNet18
@@
TinyViT We provide the tables with RA, FA, TA, RR, FR, TR, RetDev, Indisc, T-MIA and RTE for MNIST (Table XII), FashionMNIST (Table XIII), CIFAR-10 (Table XIV) and CIFAR-100 (Table XV) using the TinyViT model.
TABLE XII: MNIST - TinyViT
@@
TABLE XIII: FashionMNIST - TinyViT
@@
TABLE XIV: CIFAR-10 - TinyViT
@@
TABLE XV: CIFAR-100 - TinyViT
@@
Appendix B Per architectures rankings
Here, we present the rankings across datasets for ResNet18 (Table XVI) and TinyVit (Table XVII). We note that some methods, such as RNI or NG+, are less efficient regarding incernibility on the ViT architectures. However, methods such as SCRUB are less efficient regarding Retention Deviation on the ViT architecture.
TABLE XVI: Ranking on ResNet